9 b) For the given distribution: (5) 

 ; ,2,1,0 ,

3

1

3

2

( )  = K











P X = x = x

x

 find moment generating function, mean and 

variance of X.


1
Expert's answer
2021-03-30T07:30:20-0400

The pmf of random variable X 

"p(x)=\dfrac{e^{−λ}λ^x}{x!} ,for x=0,1,2,…."

 

So moment generating function is given by-


"M_X(t)=E[e^{tX}]=∑_{x=0}^∞e^{tx}⋅\dfrac{e^{−λ}λ^x}{x!}=e^{−λ}∑_{x=0}^∞\dfrac{(e^{tλ})^x}{x!}=e^{−λ}(e^{e^{t}λ})=e^{λ(e^t−1)}"

.

 


Now we take the first and second derivatives of  "M_X(t) ." Remember we are differentiating with respect to t :


"M'_X(t)=\dfrac{d}{dt}[e^{λ(e^t−1)}]=λe^te^{λ(e^t−1)}"

 

"M''X(t)=\dfrac{d}{dt}[λe^te^{λ(e^t−1)}]=λe^te^{λ(e^t−1)}+λ^2e^{2t}e^{λ(e^t−1)}"


Next we evaluate the derivatives at t=0 to find the first and second moments of X :


"E[X]=M'_X(0)"


"E[X^2]=M''_X(0)=\u03bbe^0e^{\u03bb(e^0\u22121)}=\u03bb=\u03bbe^0e^{\u03bb(e^0\u22121)}+\u03bb^2e^0e^{\u03bb(e^0\u22121)}=\u03bb+\u03bb^2"

 


Finally, in order to find the variance, we use the alternate formula:


"Var(X)=E[X^2]\u2212(E[X])^2=\u03bb+\u03bb^2\u2212\u03bb^2=\u03bb."